OpenStudy (anonymous):

A colony of bacteria is growing exponentially according to the function below, where t is in hours. How many hours will it take for the population to grow to 200. Round to nearest nteger B(t)=4*e^0.8t

OpenStudy (anonymous):

why are u reposting this?

OpenStudy (mertsj):

Replace B(t) with 200 and solve for t

OpenStudy (mertsj):

\[200=4e ^{.8t}\] \[50=e ^{.8t}\] \[\ln 50=\ln e ^{.8t}\] \[\ln 50=.8t \ln e\] \[3.912=.8t\] 4.89=t